OpenStudy (anonymous):

Which inequality represents the problem? A swan can dive below sea level at a rate of 3 feet per second. How many seconds will it take the swan to reach a depth of at least –32 feet?

OpenStudy (anonymous):

the answer is 10.67

OpenStudy (anonymous):

A. t+3<32 B. 3t_> 32 C. 3 2t _>3 D. t +32>3

OpenStudy (anonymous):

I think it is B) 3t_>32

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Which equation represents the problem? Pauline can spend $36 on bus fare each week. She needs to ride the bus 5 days a week. How much can she spend each day on bus fare? A. 36 + b = 5 B. 5 b = 36 C. 5 ÷ b = 36 D. 36 – b = 5

OpenStudy (anonymous):

@unkownbeast911

OpenStudy (anonymous):

i know it is not B)

OpenStudy (anonymous):

i think it is D)